EXOMOL: MOLECULAR LINE LISTS FOR ASTROPHYSICAL APPLICATIONS. A THEORETICAL LINE LIST FOR NICKEL HYDRIDE.

S. N. YURCHENKO, \underline{L. LODI}, A. KERRIDGE and J. TENNYSON, University College London, Department of Physics and Astronomy, London WC1E 6BT, UK.

Exomol (www.exomol.com) is a database of molecular line lists which can be used for spectral characterisation and simulation of astrophysical enviroments such as exoplanets, brown dwarfs, cool stars and sunspots^{a}. New line lists for about 30 small molecules of astrophysical interest which currently lack a complete spectroscopic coverage are being generated. The list includes diatomics (e.g., C_{2}, O_{2}, AlO), triatomics (e.g., H_{2}S, C_{3}, SO_{2}), tetratomics (e.g., PH_{3}, HOOH, H_{2}CO) and a few larger molecules (most notably CH_{4} and HNO_{3}). We report progress on a new theoretical line list for nickel hydride NiH. The spectra of transition-metal hydrides such as NiH are very complicated due to the large-number of low-lying electronic states, to the importance of correlation, relativistic and spin-orbit effects and of the various couplings between angular momenta. In our study potential energy curves and the relevant couplings were computed ab initio and the corresponding coupled-surface ro-vibronic problem was solved using an expansion in Hund’s case (a) wave functions. Potential curves and couplings were then refined semi-empirically using the available experimental spectroscopic data ^{b}.

aJ. Tennyson and S. N. Yurchenko, Mon. Not. R. Astron. Soc. (submitted)

bVallon R., S.H. Ashworth, P. Crozet, R.W. Field, D. Forthomme, H. Harker, C. Richard and A.J. Ross, J. Chem. Phys. A 113, 13159-13166 (2009)